fathri:

Find the slope of the line. Question options: There are a total of 21 students in the choir and the drama club. There are 7 more students in the choir than the drama club. Write and solve a system of equations to represent this solution. Let x represent the number of students in the choir, and let y represent the number of students in the drama club. Interpret the solution.   Question options: (7, 14); There are 14 students in the choir and 7 students in the drama club. (7, 14); There are 7 students in the choir and 14 students in the drama club. (14, 7); There are 14 students in the choir and 7 students in the drama club. (14, 7); There are 7 students in the choir and 14 students in the drama club.

fathri:

hello someone help me

jhonyy9:

x number of students in the choir y number of students in the drama club there are total 21 students so x+y=21 there are 7 more students in the choir than the drama club so what mean that x+7 =y x+y=21 x+7=y solve this system for x and y hope helped

fathri:

thanks

fathri:

The amount a spring will stretch varies directly with the amount of weight attached to the spring. If a spring stretches 2.1 inches when 80 pounds is attached, how far will it stretch when 40 pounds is attached?

fathri:

There are a total of 21 students in the choir and the drama club. There are 7 more students in the choir than the drama club. Write and solve a system of equations to represent this solution. Let x represent the number of students in the choir, and let y represent the number of students in the drama club. Interpret the solution.

Mercury:

next time please make sure you are only posting 1 question per post I will be addressing this one: The amount a spring will stretch varies directly with the amount of weight attached to the spring. If a spring stretches 2.1 inches when 80 pounds is attached, how far will it stretch when 40 pounds is attached? for a direct variation, y = kx where k is a constant. in this case, y is the stretch (dependent variable) and x is the weight (independent variable). in the first scenario, y = 2.1 inch and x = 80 pounds, plug in y and x to solve for k. then, in the second scenario, you have your k and x value so you can solve for y.